# Future targets for light gauge bosons from cosmic strings

### v1

@@ III.1 GW emission
The expression above predicts that, in the radiation-dominated cosmology, the GW abundance at high frequencies is constant and depends on the string tension as $\Omega_{\rm GW}(f)\propto\sqrt{\mu}$ Blanco-Pillado et al. (2014). At the opposite end, for low frequencies we find
@@
The largest emission time of GWs $t_{m}$ in Eq. (45) coincides with $t_{0}$ as long as their frequencies are sufficiently high to satisfy the condition $t_{i}<t_{\rm eq}$, ensuring that the loop generates during RD. In this context, the integration of Eq. (45) at low frequencies leads to the expression
@@
As a test, we reproduce Figure 13 from Ref. Blanco-Pillado et al. (2014) for string tensions $G\mu=10^{-7}$, $G\mu=10^{-8}$, and $G\mu=10^{-9}$. The results are shown in Fig. 1 for the GWs produced by loops formed and emitting during RD (red dashed lines), formed during RD and emitting during MD (red dotted lines), formed and emitting during MD (blue dotted lines), and the overall results (black solid lines). We find a $\mathcal{O}(1)$ discrepancy between the analytical expression in Eqs. (43) and (44) and the numerical result, which arises from the definition of $t_{\rm eq}$.
@@
III.2 Including gauge boson emission
We extend the discussion of GW emission from a cosmic string network in Sec. II.2 to include the impact of gauge boson dynamics. Since we are considering gauge boson masses significantly above $1/t_{\rm eq}\sim 10^{-27}\,$eV, their additional emission primarily affects the spectrum at high frequencies, which corresponds to loops formed during RD.
@@ III.2 Including gauge boson emission
For $k=1$, the expression in Eq. (55) predicts the behaviour $\Omega_{\rm GW}\propto f^{-1}$ at high frequencies. However, the inclusion of higher modes modifies the frequency scaling of the GW abundance to $\Omega_{\rm GW}\propto f^{-1/3}$ Blasi et al. (2020). We have verified that the numerical results obtained through Eq. (18) exhibit this behavior at high frequencies in the limit of a large number of modes. This same behaviour can also be derived analytically as follows: to recover the correct behavior at high frequencies, we approximate the summation over $k$ with an integral and take the limit at large $k$. The resulting expression is independent of the mode and reads
@@
This expression matches the findings in Ref. Blasi et al. (2020) up to a numerical prefactor.
@@ IV Discussion
We assess the model against the forecast reach of various GW measurements and the bounds obtained from cosmological considerations. Fig. 3 shows the results in terms of the model parameters: the gauge boson mass $m_{A^{\prime}}$ (horizontal axis) and the energy scale $v$ (vertical axis). The parameter space is constrained in the upper right portion by the requirement that the contribution from cold gauge bosons produced via the misalignment mechanism does not overclose the Universe Kitajima and Nakayama (2023b), see also Ref. Long and Wang (2019). This constraint is derived by enforcing $\Omega_{\rm DM}h^{2}\lesssim 0.12$ in Eq. (32) (gray shaded area). Another upper bound, not shown here, arises from requiring that the contribution of relativistic gauge bosons to the effective numbers of neutrinos does not exceed what is allowed by CMB measurements, see Eq. (36). For the range of masses considered here, this bound is not relevant, though it would become significant for lighter gauge bosons.
@@
We perform a scan over the parameter space to determine whether, for a given set of values $(m_{A^{\prime}},v)$, the corresponding GW strain falls within the forecasted reach of a given GW interferometer. We include the reach of LIGO-Virgo-Kagra (LVK) during run 2 (O2) and run 5 (O5) Abbott et al. (2009, 2018); Aasi et al. (2015); Abbott et al. (2016), represented by the green solid and blue dashed lines.
@@
We also consider the forecasts from other GW experiments. The Einstein Telescope (ET) is a proposed ground-based GW observatory expected to operate in the Hz-kHz frequency bands, with a peak strain sensitivity of $h\sim\mathcal{O}(10^{-25}){\rm\,Hz^{-1/2}}$ Punturo et al. (2010); Sathyaprakash et al. (2012); Maggiore et al. (2020). Since ET will operate within the frequency bands covered by LVK interferometers but with improved sensitivity, the forecasted region for ET shows a similar shape at lower values of $v$. The region that can be explored with ground-based interferometry weakens at low masses due to the GW spectrum cutting off at light masses, corresponding to the frequency cutoff in Eq. (58).
@@
At lower frequencies, we account for the forecast reach of the upcoming Laser Interferometer Space Antenna (LISA) satellite Amaro-Seoane et al. (2012); Barausse et al. (2020); Auclair et al. (2023); Blanco-Pillado et al. (2024), which has been recently funded by the European Space Agency (ESA) and is planned for launch in the 2030s. This is represented by the area bounded by the dashed purple line in Fig. 3. A proposed successor to LISA by ESA is the Big Bang Observer (BBO), which aims to probe the mHz to kHz frequency range, peaking at $f_{\rm GW}\sim 1\,$Hz with a strain sensitivity of $|h|\sim 10^{-24}{\rm\,Hz^{-1/2}}$ Yagi and Seto (2011). BBO will extend the search for cosmic strings from a phase transition well below the current limits, reaching down to the symmetry-breaking scale $v\lesssim 10^{8}\,$eV.
@@
At even lower frequencies, the proposed $\mu$Ares experiment aims to cover the GW bands in the $\mu$Hz to mHz frequency range Sesana et al. (2021). An overview of the expected sensitivities for GW interferometers can be found in Refs. Schmitz (2021, 2020). The results of the scan for all the considered interferometer forecasts are shown in Fig. 3, with specific curves for LVK O2 (green), LVK O5 (blue), ET (pink), LISA (purple), $\mu$Ares (red), and BBO (orange). Besides LVK which has already concluded its run O2 and it is represented by a solid line, all expected forecast reaches are shown with dashed lines. Since BBO, LISA and $\mu$Ares operate below the mHz range, these interferometers are unable to probe the pivot frequency which scales with the string energy density as $f_{*}\propto(G\mu)^{-1}$. Consequently, they cannot determine the nature of the string network (global versus gauge). The results for LISA align with those obtained for $\alpha=0.1$ in Ref. Auclair et al. (2020).
@@
The scaling at extremely low frequencies, corresponding to the nHz range, could potentially contribute to the detection of the GWB through pulsar timing analysis. Earlier work constrains the string tension as $G\mu\lesssim 8\times 10^{-10}$ using data from the European Pulsar Timing Array (EPTA) Lentati et al. (2015), $G\mu\lesssim 5.3\times 10^{-11}$ using data from North American Nanohertz Observatory for Gravitational Waves (NANOGrav) Arzoumanian et al. (2018), or $G\mu\lesssim 5.1\times 10^{-10}$ using data from Parkes Pulsar Timing Array (PPTA) Chen et al. (2022). The recent detection of the GW spectrum at such low frequencies suggests a string tension in the range $-11\lesssim\log_{10}G\mu\lesssim-9.5$, with the exact values depending on the specific distribution of loops in the string network Antoniadis et al. (2024); Afzal et al. (2023). This is illustrated by the cyan shading in Fig. 3. Also shown is the constraint from BH superradiance Arvanitaki et al. (2010); Arvanitaki and Dubovsky (2011) derived from the observation of highly spinning black holes (red dashed area), which restricts the mass of the gauge boson as $10^{-13}{\rm\,eV}\lesssim m_{A^{\prime}}\lesssim 3\times 10^{-12}{\rm\,eV}$ Cardoso et al. (2018).
@@
As anticipated in the discussion around Fig. 3, we have incorporated additional bounds from complementary sources. Constraints on the number of effective degrees of freedom $N_{\rm eff}$ have been used to place an upper limit on the SSB scale $v$ due to the production of relativistic gauge bosons. These measurements also constrain the abundance of the stochastic GW signal at various frequencies, with stricter bounds applied at higher frequencies. This constraint impacts the amount of GWs produced by cosmic strings of cosmological origin, as these GWs would be present during BBN and at CMB decoupling. At BBN, this results in Caprini and Figueroa (2018)
@@
where $\rho_{\rm GW}$ refers to the portion of the GWB that is present before BBN occurs. Following Eq. (27), the energy density in GWs is Servant and Simakachorn (2024)
@@
We adopt the recent results for $\Delta N_{\rm eff}$ from the Planck satellite Aghanim et al. (2020),
@@ V Conclusions
For all of these proposed or planned interferometers, we perform a scan over the parameters $(m_{A^{\prime}},v)$ to decide which region of the parameter space could be covered with future missions. The results of this analysis are summarized in Fig. 3, following the comparisons exemplified in Fig. 4. Future work on this topic will revolve around the understanding of the GW spectrum in scenarios where the Abelian Higgs model and its extensions are relevant. This includes, but is not limited to, exploring the implications of the $B\textrm{--}L$ gauge symmetry, which plays a crucial role in many grand unified theories Wilczek and Zee (1979) and physics beyond the Standard Model Escudero et al. (2018); Lin et al. (2022).

### v2

@@ 3.1 GW emission
The expression above predicts that, in the radiation-dominated cosmology, the GW abundance at high frequencies is constant and depends on the string tension as $\Omega_{\rm GW}(f)\propto\sqrt{\mu}$ [71]. At the opposite end, for low frequencies we find
@@
In Eq. (45), the largest emission time $t_{m}$ of GWs coincides with $t_{0}$, provided the frequencies are high enough to satisfy the condition $t_{i}<t_{\rm eq}$, ensuring that the loop forms during RD. In this context, the integration of Eq. (45) at low frequencies yields the expression
@@
As a test, we reproduce Figure 13 from Ref. [71] for string tensions $G\mu=10^{-7}$, $G\mu=10^{-8}$, and $G\mu=10^{-9}$. The results are shown in Fig. 1 for the GWs produced by loops formed and emitting during RD (red dashed lines), formed during RD and emitting during MD (red dotted lines), formed and emitting during MD (blue dotted lines), and the overall results (black solid lines). We find a $\mathcal{O}(1)$ discrepancy between the analytical expression in Eqs. (43) and (44) and the numerical result, which arises from the definition of $t_{\rm eq}$.
@@
3.2 Including gauge boson emission
We extend the discussion of GW emission from a cosmic string network in Sec. 2.2 to include the impact of gauge boson dynamics. Since we are considering gauge boson masses significantly above $1/t_{\rm eq}\sim 10^{-27}\,$eV, their additional emission primarily affects the spectrum at high frequencies, which corresponds to loops formed during RD.
@@ 3.2 Including gauge boson emission
For $k=1$, the expression in Eq. (55) predicts the behaviour $\Omega_{\rm GW}\propto f^{-1}$ at high frequencies. However, the inclusion of higher modes modifies the frequency scaling of the GW abundance to $\Omega_{\rm GW}\propto f^{-1/3}$ [96, 37, 97]. We have verified that the numerical results obtained through Eq. (18) exhibit this behavior at high frequencies in the limit of a large number of modes. This same behaviour can also be derived analytically as follows: to recover the correct behavior at high frequencies, we approximate the summation over $k$ with an integral and take the limit at large $k$. The resulting expression is independent of the mode and reads
@@
This expression matches the findings in Ref. [96] up to a numerical prefactor.
@@ 4 Discussion
We assess the model against the forecast reach of various GW measurements and the bounds obtained from cosmological considerations. Fig. 3 shows the results in terms of the model parameters: the gauge boson mass $m_{A^{\prime}}$ (horizontal axis) and the energy scale $v$ (vertical axis). The parameter space is constrained in the upper right portion by the requirement that the contribution from cold gauge bosons produced via the misalignment mechanism does not overclose the Universe [57, 59]. This constraint is derived by enforcing $\Omega_{\rm DM}h^{2}\lesssim 0.12$ in Eq. (32) (gray shaded area). Note that the result shown in Fig. 3 below includes the estimate for DM abundance reported in Ref. [58], which differs from the analysis in Ref. [57] by a factor $\sim\mathcal{O}(1)$. This discrepancy likely arises from unresolved differences between the simulations, particularly concerning the abundance of transverse vector bosons present when the network dissipates. We will not delve further into this discussion here, as our goal is to highlight the potential target for future experiments. Another upper bound, not shown here, comes from the requirement that the contribution of relativistic gauge bosons to the effective number of neutrinos does not exceed the limit allowed by CMB measurements, see Eq. (36). For the mass range considered here, this bound is not relevant, though it would become significant for lighter gauge bosons.
@@
We perform a scan over the parameter space to determine whether, for a given set of values $(m_{A^{\prime}},v)$, the corresponding GW strain falls within the forecasted reach of a given GW interferometer. We include the reach of LIGO-Virgo-Kagra (LVK) during run 2 (O2) and run 5 (O5) [98, 99, 100, 101], represented by the green solid and blue dashed lines.
@@
We also consider the forecasts from other GW experiments. The Einstein Telescope (ET) is a proposed ground-based GW observatory expected to operate in the Hz-kHz frequency bands, with a peak strain sensitivity of $h\sim\mathcal{O}(10^{-25}){\rm\,Hz^{-1/2}}$ [102, 103, 104]. Since ET will operate within the frequency bands covered by LVK interferometers but with improved sensitivity, the forecasted region for ET shows a similar shape at lower values of $v$. The region that can be explored with ground-based interferometry weakens at low masses due to the GW spectrum cutting off at light masses, corresponding to the frequency cutoff in Eq. (58).
@@
At lower frequencies, we account for the forecast reach of the upcoming Laser Interferometer Space Antenna (LISA) satellite [105, 106, 107, 108], which has been recently funded by the European Space Agency (ESA) and is planned for launch in the 2030s. This is represented by the area bounded by the dashed purple line in Fig. 3. A proposed successor to LISA by ESA is the Big Bang Observer (BBO), which aims to probe the mHz to kHz frequency range, peaking at $f_{\rm GW}\sim 1\,$Hz with a strain sensitivity of $|h|\sim 10^{-24}{\rm\,Hz^{-1/2}}$ [109]. BBO will extend the search for cosmic strings from a phase transition well below the current limits, reaching down to the symmetry-breaking scale $v\lesssim 10^{8}\,$eV.
@@
At even lower frequencies, the proposed $\mu$Ares experiment aims to cover the GW bands in the $\mu$Hz to mHz frequency range [110]. An overview of the expected sensitivities for GW interferometers can be found in Refs. [111, 112]. The results of the scan for all the considered interferometer forecasts are shown in Fig. 3, with specific curves for LVK O2 (green), LVK O5 (blue), ET (pink), LISA (purple), $\mu$Ares (red), and BBO (orange). Besides LVK which has already concluded its run O2 and it is represented by a solid line, all expected forecast reaches are shown with dashed lines. Since BBO, LISA and $\mu$Ares operate below the mHz range, these interferometers are unable to probe the pivot frequency which scales with the string energy density as $f_{*}\propto(G\mu)^{-1}$. Consequently, they cannot determine the nature of the string network (global versus gauge). The results for LISA align with those obtained for $\alpha=0.1$ in Ref. [38].
@@
The scaling at extremely low frequencies, corresponding to the nHz range, could potentially contribute to the detection of the GWB through pulsar timing analysis. Earlier work constrains the string tension as $G\mu\lesssim 8\times 10^{-10}$ using data from the European Pulsar Timing Array (EPTA) [113], $G\mu\lesssim 5.3\times 10^{-11}$ using data from North American Nanohertz Observatory for Gravitational Waves (NANOGrav) [114], or $G\mu\lesssim 5.1\times 10^{-10}$ using data from Parkes Pulsar Timing Array (PPTA) [115]. The recent detection of the GW spectrum at such low frequencies suggests a string tension in the range $-11\lesssim\log_{10}G\mu\lesssim-9.5$, with the exact values depending on the specific distribution of loops in the string network [116, 117]. This is illustrated by the cyan shading in Fig. 3. Also shown is the constraint from BH superradiance [118, 119] derived from the observation of highly spinning black holes (red dashed area), which restricts the mass of the gauge boson as $10^{-13}{\rm\,eV}\lesssim m_{A^{\prime}}\lesssim 2\times 10^{-11}{\rm\,eV}$ [120, 121, 122]. We restrict ourselves to the results obtained for masses $m_{A^{\prime}}\gtrsim 10^{-15}$ eV for a few reasons. Firstly, lighter bosons in the mass range $10^{-19}{\rm\,eV}\lesssim m_{A^{\prime}}\lesssim 10^{-16}{\rm\,eV}$ are disfavored by the observation of spinning supermassive black holes at various mass ranges above $10^{6}\,M_{\odot}$ [123, 122]. Heavier black holes with masses in the range $\gtrsim 10^{9}\,M_{\odot}$ could even probe the mass scale below $10^{-21}$ eV [124, 125], although such massive BHs lack precise measurements of their spins. Besides, various cosmological measurements disfavor a large portion of the DM being in the form of a light boson, such as the Lyman-$\alpha$ absorption lines caused by neutral hydrogen in the intergalactic medium and constraining to $m_{A^{\prime}}\gtrsim 10^{-21}$ eV [126, 127, 128, 129].
@@
We do not expect the gauge boson mass to be ultralight in models that connect the dark photon to the Standard Model. The mass of the light boson is generally related to the vacuum energy scale $v$ as $m_{A’}\sim gv$, where $g$ is the gauge boson coupling. In models based on gauged symmetries, such as baryon number conservation, the gauge boson is expected to couple with neutrinos, implying that $v$ is linked to beyond the Standard Model physics at a high energy scale. The gauge boson coupling is also loosely constrained by the weak gravity conjecture [130], which requires $g\gtrsim m_{\nu}/m_{\rm Pl}\sim 10^{-30}$ for a neutrino mass $m_{\nu}\sim 10^{-2}$ eV. For example, setting $v\gtrsim 10^{8}$ GeV, which is the lowest energy considered in Fig. 3, leads to $m_{A’}\gtrsim 10^{-13}$ eV. Although model building allows for theories with a gauge boson mass spanning various orders of magnitude [131, 132, 133, 134, 135], we do not consider ultralight bosons here.
As anticipated in the discussion around Fig. 3, we have incorporated additional bounds from complementary sources. Constraints on the number of effective degrees of freedom $N_{\rm eff}$ have been used to place an upper limit on the SSB scale $v$ due to the production of relativistic gauge bosons. These measurements also constrain the abundance of the stochastic GW signal at various frequencies, with stricter bounds applied at higher frequencies. This constraint impacts the amount of GWs produced by cosmic strings of cosmological origin, as these GWs would be present during BBN and at CMB decoupling. At BBN, this results in [28]
@@
where $\rho_{\rm GW}$ refers to the portion of the GWB that is present before BBN occurs. Following Eq. (27), the energy density in GWs is [43]
@@
We adopt the recent results for $\Delta N_{\rm eff}$ from the Planck satellite [95],
@@ 5 Conclusions
For all of these proposed or planned interferometers, we perform a scan over the parameters $(m_{A^{\prime}},v)$ to decide which region of the parameter space could be covered with future missions. The results of this analysis are summarized in Fig. 3, following the comparisons exemplified in Fig. 4. Future work on this topic will revolve around the understanding of the GW spectrum in scenarios where the Abelian Higgs model and its extensions are relevant. This includes, but is not limited to, exploring the implications of the $B\textrm{--}L$ gauge symmetry, which plays a crucial role in many grand unified theories [136] and physics beyond the Standard Model [137, 138].